OpenStudy (anonymous):

1.Solve the quadratic equation x^(2) + 12x + 20 = 0 by completing the square. 2.Use the quadratic formula to solve 5x2 − 2x − 3 = 0

OpenStudy (anonymous):

For the first \[\{x= -10\},\{x= -2\}\} \] For the second \[\left\{\left\{x= -\frac{3}{5}\right\},\{x= 1\}\right\} \]

OpenStudy (anonymous):

2. 2 + square root of -54/10 and 2- square root of -54/10 That's what i got at least :)

OpenStudy (anonymous):

1. x=-2,-10